OpenStudy (anonymous):

Solve P = 2(l + w) for l. What are the missing values in the table? P w l 14 2 ? 22 8 ?

OpenStudy (anonymous):

Use the given equation, for the first case plug P = 14 and w = 2, you'd get\[14=2(l+2)\]now solve for l\[7= (l+2)\]

OpenStudy (lopus):